Using the principle of mathematical induction, prove that

(n5/5) + (n3/3) + (7n/15) is a natural number, ∀ n ∈ N .


1
Expert's answer
2021-05-31T19:09:43-0400

Let "P(n)=\dfrac{n^{5}}{5}+\dfrac{n^{3}}{3}+\dfrac{7n}{15}" is a natural number for all "n" "\u2208N"

now putting "n=1" , we get -


"P(1)=\dfrac{1^{5}}{5}+\dfrac{1^{3}}{3}+\dfrac{7\times1}{15}=\dfrac{15}{15}=1" .


which is a natural number .


Hence , P(1) is true .


Let us prove that P(n) is true for some natural number n=k.


"\therefore" "P(k)=\dfrac{k^{5}}{5}+\dfrac{k^{3}}{3}+\dfrac{7k}{15}" is a natural number ..........(1)




Now we have to prove that "P(k+1)" is true.


"P(k+1)=" "\dfrac{(k+1)^{5}}{5}+\dfrac{(k+1)^{3}}{3}+\dfrac{7(k+1)}{15}"


"=" "\dfrac{k^{5}+5k^{4}+10k^{3}+10k^{2}+5k+1}{5}+\dfrac{k^{3}+1+3k^{2}+3k}{3}+\dfrac{7k+7}{15}"


"=\dfrac{k^{5}}{5}+\dfrac{k^{3}}{3}+\dfrac{7k}{15}+" "\dfrac{5k^{4}+10k^{3}+10k^{2}+5k+1}{5}+" "\dfrac{1+3k^{2}+3k}{3}+\dfrac{7}{15}"


"=" "\dfrac{k^{5}}{5}+\dfrac{k^{3}}{3}+\dfrac{7k}{15} \ +k^{4}+2k^{3}+2k^{2}+k+k^{2}+k+\dfrac{1}{5}+\dfrac{1}{3}+\dfrac{7}{15}"


"=" "\dfrac{k^{5}}{5}+\dfrac{k^{3}}{3}+\dfrac{7k}{15}+k^{4}+2k^{3}+3k^{2}+2k+1"



which is a natural number .....by using .....(1)



Hence , "P(k+1)" is true whenever P(k) is true.


so by principal of mathematical induction


"P(n)" is true for any natural number n .